# The thermodynamic structure and large-scale structure filament in MACS J0717.5+3745

### v2

@@ Appendix B Instrumental Cross-calibration
Regarding measured temperatures, previous studies have shown a 5-10% difference on the effective temperature between the MOS and pn detectors (Schellenberger et al., 2015). Chandra systematically has shown higher temperatures than XMM-Newton in the broad band, yet in another more recent study using Chandra and NuSTAR data, Wallbank et al. (2022) showed a 10.5% and a 15.7% temperature discrepancy between the instruments using the broad band (0.6 - 9.0 keV) and hard band (3.0 - 10.0 keV), respectively. An even more recent cross-calibration talk from the 2023 IACHEC meeting [^6] https://iachec.org/wp-content/presentations/2023/WikIACHEC2023.pdf elaborated on this study further and discusses a complex relationship between the cross-calibrations between Chandra and XMM-Newton, using NuSTAR as a reference. This talk mentions that the temperature discrepancies between instruments increase with respect to measured temperatures, also noting a $\sim$20flux excess for XMM-Newton in the soft band. Furthermore, there is evidence that XMM-Newton agrees better with NuSTAR for hot clusters, while Chandra agrees better for cooler clusters.
@@
We performed several independent tests to explore the cross-calibrations between the various instruments as part of the verification of our joint fitting methodology using the instrumental background models. We found a consistent, 12-17% cross-calibration related uncertainty between Chandra and XMM-Newton between the three different observations of MACS J0717.5+3745 and a 1-3% uncertainty between the MOS and pn detectors. To do this, we fit a circular region from the high S/N regime of MACS J0717.5+3745, where there is an extremely hot ($\gtrsim 20$ keV) plasma emission, using a collisional ionisation equilibrium (apec) plasma model multiplied by a constant. We further explored the effect that the scaling of the instrumental backgrounds have on the fit results by scaling the Chandra backgrounds by the number of counts and also by scaling by the count-rate in the 10-12 keV band. We found that the temperatures were consistent between Chandra and XMM-Newton when the background was scaled by the ratio of the observation to background high-energy particle count rates, but that the normalizations were off by around 18%. However, when the background is instead scaled by the ratio of the high-energy number of counts rather than by exposure time, the normalizations were within 5%, yet the temperatures were then off by around 22%.
Since we are jointly fitting all instruments from Chandra and XMM-Newton, the systematic uncertainties are accounted for in the averaged model fit between all instruments. Because MACS J0717.5+3745 is an exceptionally hot cluster, this means that in the high S/N regime where there are many counts, a joint fit will be more correctly weighted by the larger collecting area of XMM-Newton, while near the cooler cluster outskirts, this bias will be mitigated by the inclusion of the Chandra data. Despite our measured discrepancies between the different instruments, we found that our uncertainties fell within the previously reported limits from different IACHEC studies, so our new instrumental background modelling approach does not seem to be introducing any new biases to the measurements that were not already previously known or reported.

### v3

@@ Appendix B Instrumental Cross-calibration
Regarding measured temperatures, previous studies have shown a 5-10% difference on the effective temperature between the MOS and pn detectors (Schellenberger et al., 2015). Chandra systematically has shown higher temperatures than XMM-Newton in the broad band, yet in another more recent study using Chandra and NuSTAR data, Wallbank et al. (2022) showed a 10.5% and a 15.7% temperature discrepancy between the instruments using the broad band (0.6 - 9.0 keV) and hard band (3.0 - 10.0 keV), respectively. An even more recent cross-calibration talk from the 2023 IACHEC meeting [^6] https://iachec.org/wp-content/presentations/2023/WikIACHEC2023.pdf elaborated on this study further and discusses a complex relationship between the cross-calibrations between Chandra and XMM-Newton, using NuSTAR as a reference. This talk mentions that the temperature discrepancies between instruments increase with respect to measured temperatures, also noting a $\sim$20% flux excess for XMM-Newton in the soft band. Furthermore, there is evidence that XMM-Newton agrees better with NuSTAR for hot clusters, while Chandra agrees better for cooler clusters.
@@
We performed several independent tests to explore the cross-calibrations between the various instruments as part of the verification of our joint fitting methodology using the instrumental background models. We found a consistent, 12-17% cross-calibration related uncertainty between Chandra and XMM-Newton between the three different observations of MACS J0717.5+3745 and a 1-3% uncertainty between the MOS and pn detectors. To do this, we fit a circular region from the high S/N regime of MACS J0717.5+3745, where there is an extremely hot ($\gtrsim 20$ keV) plasma emission, using a collisional ionisation equilibrium (apec) plasma model multiplied by a constant. We further explored the effect that the scaling of the instrumental backgrounds has on the fit results by scaling the Chandra backgrounds by the number of counts and also by scaling by the count-rate in the 10-12 keV band. We found that the temperatures were consistent between Chandra and XMM-Newton when the background was scaled by the ratio of the observation to background high-energy particle count rates, but that the normalizations were off by around 18%. However, when the background is instead scaled by the ratio of the high-energy number of counts rather than by exposure time, the normalizations were within 5%, yet the temperatures were then off by around 22%. Despite our measured discrepancies between the different instruments, we found that our uncertainties fell within the previously reported limits from different IACHEC studies, so our new instrumental background modelling approach does not seem to be introducing any new biases to the measurements that were not already previously known or reported.
Fig. 13 shows the statistical errors of the various instruments as a function of temperature using our joint fitting method, by extracting spectra from various regions with different temperatures. For comparison, the systematic uncertainties for temperatures of up to $\sim 10$ keV, are estimated to be of the order of 10–15% Nevalainen et al. (2010), while for higher temperatures, based on the studies of the Bullet cluster, they are around 20% (Markevitch, 2006; Wik et al., 2014).
Figure 13: Plot of the 1 sigma error as a function of temperature for Chandra, XMM-Newton, and the combined fitting procedure in units of keV and as percents.
@@ Appendix C MUSTANG2 Full FOV Image
Figure 15: A realization of the MUSTANG2 noise map across the FOV shown in Fig. 14. The FOV corresponds to the white box region shown throughout the paper, while the green box, and contours are the same as described in Fig. 11.
Appendix D 3D Line of Sight Adjusted Maps and Projection Effects
The magnitude and spatial distribution of the geometric correction are clearly highlighted in Fig.16, showing fractional differences between corrected and original maps. Additionally, Fig.17 presents these differences expressed in terms of statistical significance ($\sigma$), based on the propagated uncertainties calculated from the original spectral fit for each independent bin. We caution the reader that, due to this being a highly disturbed, merging cluster with at least five subclusters in various stages of merger, the geometry assumed for this estimate is highly uncertain.
Figure 16: Fractional differences between the corrected thermodynamic maps and the original maps, illustrating the scale of corrections due to the assumed LoS geometry.
Figure 17: Differences between the corrected and original maps expressed in units of sigma ($\sigma$), calculated from the statistical uncertainties associated with each bin.